Will the following system be satisfied by the point(-3, 5)?
1/3x +5 = 4
-1/5y - 3x = 10

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

Sub -3 for x and 5 for y and see.

Answer by richwmiller(17219)

Plug in -3 . Do you get 5 for y?
BTW you are missing y in the first equation.
Is it 5y or 4y?
Just using the second equation
-1/5*y-3x=10
There is a question whether it is -1/5*y or (-1/5)*y
-1/5*5-3(-3)=10
-1+9=10
8=10
The point doesn't work this way.
-1/(5y)-3*(-3)=10
-1/25+9=10
It doesn't work this way either.
So the point does work in equation two so it isn't a solution for the system.
Whether or not it works in the first equation.